OpenStudy (anonymous):

What is the equation of the line that has a slope of 2/3 and a y - intercept of -1/2?

OpenStudy (anonymous):

It would be Y = 2/3X + - 1/2 Becaause the formula is Y = MX + B and m is your slope and B is your y intercept